OpenStudy (anonymous):

Int(x^3/(3x-6)) dx Show full steps

OpenStudy (anonymous):

u=3x-6 du=dx (u+6)^3/27=x^3 u2+12u+36(u+6) int1/27 u^3+18u^2+108u+216/u 1/27(u^2+18u+108+216/u)du then do each individually 1/27(1/3u^3+9u^2+108u+216ln(u)+c

OpenStudy (anonymous):

times 1/3 from the du=1/3dx

OpenStudy (anonymous):

1/81(1/3u^3+9u^2+108u+216ln(u))+C

OpenStudy (anonymous):

i could be very very wrong i didnt do any of this on paper

OpenStudy (zarkon):

I would have started with long division to get \[\frac{x^3}{3x-6}=\frac{8}{3(x-2)}+\frac{x^2}{3}+\frac{2x}{3}+\frac{4}{3}\] then integrated term by term.

OpenStudy (anonymous):

dont really understand whats going on with your u substitution bnut and i have a hard time understanding what you did there Zarkon

OpenStudy (anonymous):

ok i get now on paper 1/3( u^3/3+3u^2+12u+8ln(u)) u=x-2 x^3=(x+2)^3

OpenStudy (zarkon):

have you done polynomial division before?

OpenStudy (anonymous):

yes i have i just now figured out how you did it XD

OpenStudy (anonymous):

ok so i said u=3x-6 do du/3=dx then i isolated x=(u+6)/3 then i cubed that to get x^3

OpenStudy (anonymous):

ahh ok i see i see, gotcha

OpenStudy (zarkon):

you should end up with \[\frac{8}{3}\ln(x-2)+\frac{x^3}{9}+\frac{x^2}{3}+\frac{4}{3}x+c\]

OpenStudy (anonymous):

ahh i see thats one good way to do it zarkon, wana see bnuts sub way too once hes done with it

OpenStudy (zarkon):

ya...it is good to know how to do it both ways.

OpenStudy (anonymous):

do you wanna see a good example of it i did it on paper but zarkon ways is a lot more fluid than mine, theres a alot of room for mistake like you saw me do. \[ u=3x-6\]

OpenStudy (anonymous):

well you can if you want to i wouldnt mind seeing another example but if not its kool too bnut

OpenStudy (anonymous):

\[(u+6)^3\div3^3=x^3\] do you agree up tho here?

OpenStudy (anonymous):

yosh

OpenStudy (anonymous):

and the du/3=dx and you have all your subs in there so your integral would look like this\[1/27\int\limits_{}^{}((u+6)^3/u)du \] then youll get a bunch od u's cancelling and the last term should be a CONSTANT which is where the ln(u) comes into play. try it out

OpenStudy (anonymous):

ahhh just tried it out, neat!

OpenStudy (anonymous):

but the other way would be easier, but calculus is def on how creative you can be when doing integrals, der, or any apps good luck with the class

OpenStudy (anonymous):

yah i see what you mean bnut, thanks for your help as well XD